# ASTERIX: Module for modelling the water flow on vegetated hillslopes

### v1

@@ 5 Conclusions and final remarks
The software is designed to work on regular spatial hexagonal network. When the terrain data are provided by a GIS file, the software uses the “porting data” module [22] which transfers the information to a hexagonal raster. It can also directly generate information (topography, vegetation, etc) on hexagonal rasters for theoretical or laboratory experiments. For modelling the water flow, one can start with an initial water layer or use a rain simulator (e.g. a triangular hyetograph as in C).
To run ASTERIX, one must know the values of some environmental parameters in order to get results as close to the real phenomena as possible. For example, the friction coefficients $\alpha_{s}$ and $\alpha_{p}$ cannot be directly measured; instead, one can use some reference values from literature or they can be evaluated through some inverse methods using laboratory measurements to fit the data, see Subsection 4.4 as a sketch example. Note that the estimation of these friction parameters is influenced by the distribution of the vegetation. For example, it is more difficult to locate $\alpha_{s}$ on the flume experiments with downstream vegetation (Fig. 11, Case II) and to locate $\alpha_{p}$ when the vegetation is placed upstream (Fig. 11, Case III). Fig. 11, Case I emphasizes that one can find a large number of pairs $(\alpha_{s},\alpha_{p})$ leading to very small errors and that it is difficult to decide the correct parameters when only one experiment is used. Their best localization is obtained when using all the experiments (Fig. 11, Case IV). We can also imagine other scenarios (with variable $\theta$ or different values of the slope $m$, for example), but we did not go further because we wanted to compare the results with experimental measurements. This investigation suggest that one must cleverly choose a set of experiments with different configurations in order to determine the parameters as best as possible.
In real conditions, the terrain is heterogeneous and working only with average values for $\alpha_{s}$ and $\alpha_{p}$ can be risky and, therefore, distributed values on the entire domain should be used instead. In such cases, a possible solution is to partition the surface into quasi-homogeneous subdomains with constant friction parameters.
@@ Appendix A Details on the Numerical Scheme
$\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{*}$ $=\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{n}+\triangle t_{n}\left({\cal J}^{v}_{a\;i}((h,\boldsymbol{v})^{n})+{\cal S}_{a\;i}((h,w)^{n})\right),$ (35b)
@@
This second step has an analytic solution of the form
@@
$\alpha:=(\theta_{i}h_{i})^{n+1},\quad\beta:=\triangle t_{n}{\cal K}(h_{i}^{n+1},\theta_{i}),\quad\gamma_{a}:=(\theta_{i}h_{i}v_{a\;i})^{*}.$
@@
Appendix B Analytic Solution for Thacker Problem
@@
$W_{0}(t;\Delta,V_{0},V^{\prime}_{0}):=\frac{1}{2g}\left[A^{2}\frac{\lambda_{2}}{\lambda_{1}}\left(1-e^{2\lambda_{1}t}\right)+B^{2}\frac{\lambda_{1}}{\lambda_{2}}\left(1-e^{2\lambda_{2}t}\right)+2AB\left(1-e^{-\tau t}\right)\right]$ (46)
$\begin{split}w_{0}(t):=&w_{0}(t;\Delta_{a},\Delta_{b},u_{0},v_{0},u^{\prime}_{0},v^{\prime}_{0})\\ :=&w_{0}+W_{0}(t;\Delta_{a},u_{0},u^{\prime}_{0})+W_{0}(t;\Delta_{b},v_{0},v^{\prime}_{0}),\end{split}$ (47)
@@ Appendix C Rainfall Hyetograph: triangular model
The graph of the instantaneous rain intensity $r(t)$ over the time interval $[0,T_{d}]$ has the shape of a triangle with $T_{d}$ and $i_{a}$ as its base and height, respectively (see Fig. 15).
@@
$r(t)=\left\{\begin{array}[]{ll}i_{a}\displaystyle\frac{t}{t_{a}},&0\leq t\leq t_{a},\\ i_{a}\displaystyle\frac{1-t/T_{d}}{1-\gamma},&t_{a}<t\leq T_{d},\\ 0,&{\rm otherwise}.\end{array}\right.$

### v2

@@ 5 Conclusions and final remarks
The software is designed to work on regular spatial hexagonal network. The software provides a “porting data” module [23] for generating a hexagonal raster by using data from a rectangular raster or analytical functions. The hexagonal raster can also be used for purposes other than supporting grid for numerical scheme, e.g flow routing or cellular automata. The core of the software consists of two modules: porting data numerical integration. Both modules were designed based on two principles: as simple and as “universal” as possible. By universal we mean that the software can be used for a large class of physical environments: topography, land use, hydrographic basins, soil physics, meteorology, etc. We keep it simple in order to allow one to add new modules to cope with given specific conditions or requirements, e.g. boundary conditions varying in time and space, distributed parameters or a post-processing data module. The source code and the user manual contain many details to facilitate the modification of the existent functions or to introduce new ones. The user are free to do all he/she wishes.
A lot of work was spent to create a video interface that allows the visualization of the evolution of the hydrodynamic variables over time. This facility can be informative for hydrologist or environmental policymaker.
Another purpose of the software is to be used as a part of an inverse method designed to estimate some physical parameters of the mathematical model. In this sense, the numerical integration module has been planned to run independently from the other modules. The Subsection 4.3 presents an example of an inverse method to estimate the coefficients $\alpha_{s}$ and $\alpha_{p}$ based on a laboratory experiment. Unfortunately, the problem of parameter estimation by an inverse method is ill-posed in the sense that very poor localization of minimal point of the cost function or amplification of the error in data are reached. For example, we found that both parameters $\alpha_{s}$ and $\alpha_{p}$ are very poorly located when the functional cost is built up on a single experiment, but they are very well located when one use more and different experiments (Case IV).
@@
We can also imagine other scenarios (with variable $\theta$ or different values of the slope $m$, for example), but we did not go further because we wanted to remain inside the frame of experimental measurements. This investigation suggests that one must cleverly choose a set of experiments with different configurations in order to determine the parameters as best as possible.
In real conditions, the terrain is heterogeneous and working only with a single value for $\alpha_{s}$ and a single value for $\alpha_{p}$ can be risky and therefore, distributed values on the entire domain should be used instead. In such cases, a possible solution is to partition the surface into quasi-homogeneous subdomains with constant friction parameters.
The estimation of the porosity function is an important issue from a real application point of view. For the case of well structured cover plant (agricultural land, forest or laboratory experiment), one can easily estimate the area of domain occupied by the plant stems, but for other cases this estimation is very difficult. A possible solution to this problem is to combine the terrain pointwise determination with LiDAR measurements or aerial photography. Another solution is to set up an inverse method to find the porosity function.
@@
CRediT authorship contribution statement
All three authors contributed equally to this work: Conceptualization, Methodology, Software, Validation, Formal analysis, Data Curation, Writing - Original Draft.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
[…]
@@ Appendix A Details on the Numerical Scheme
$\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{*}$ $=\sigma_{i}(\theta_{i}h_{i}v_{a\;i})^{n}+\triangle t_{n}{\cal J}^{v}_{a\;i}((h,\boldsymbol{v})^{n})+\triangle t_{n}{\cal S}_{a\;i}((h,w)^{n}),$ (31b)
@@
Solving (16) is accomplished in three steps:
  1. 1.
    Set $\triangle t_{n}$ subject to the restriction
    $\triangle t_{n}<\tau_{n},$
    where
    $\tau_{n}=CFL\displaystyle\frac{\phi_{\rm min}}{c^{n}_{\rm max}},$ (33)
    with
    $\begin{array}[]{c}c_{i}=|\boldsymbol{v}|_{i}+\sqrt{gh_{i}},\quad c_{\rm max}=\max\limits_{i}\{c_{i}\},\\ \phi_{\rm min}=\min\limits_{i}\left\{\displaystyle\frac{\sigma_{i}}{\sum\limits_{j\in{\cal N}(i)}l_{(i,j)}}\right\}.\end{array}$ (34)
  2. 2.
    Calculate $h_{i}^{n+1}$ using the mass balance relation of (16).
  3. 3.
The time step limitation is imposed by the hyperbolic character of the shallow-water equations and by the positivity-preserving requirement of the water depth. $CFL$ is a number between 00 and $1$ (the Courant-Friedrichs-Lewy condition). The reader is referred to Appendix A for extended details on the last two steps.
When the software is used to study the Riemann Problem or models that develop shock type discontinuity, it is recommended [35, 50] to use an augmented linear momentum flux ${\cal J}^{\boldsymbol{v}}_{a\,i}$ instead of ${\cal J}_{a\,i}(h,\boldsymbol{v})$:
This update is based on the artificial viscosity
where
$c_{(i,j)}=\max\{c_{i},c_{j}\},\qquad\mu_{(i,j)}=\displaystyle\frac{2\theta_{i}h_{i}\theta_{j}h_{j}}{\theta_{i}h_{i}+\theta_{j}h_{j}}.$ (37)
This last step has an analytic solution of the form
@@
$\alpha:=(\theta_{i}h_{i})^{n+1},\qquad\beta:=\triangle t_{n}{\cal K}(h_{i}^{n+1},\theta_{i}),\qquad\gamma_{a}:=(\theta_{i}h_{i}v_{a\;i})^{*}.$
@@
Appendix B Analytic Solution for Thacker’s Problem
@@
$W_{0}(t;\Delta,V_{0},V^{\prime}_{0}):=\frac{1}{2g}\left[A^{2}\frac{\lambda_{2}}{\lambda_{1}}\left(1-e^{2\lambda_{1}t}\right)+B^{2}\frac{\lambda_{1}}{\lambda_{2}}\left(1-e^{2\lambda_{2}t}\right)+2AB\left(1-e^{-\tau t}\right)\right],$ (47)
$w_{0}(t):=w_{0}(t;\Delta_{a},\Delta_{b},u_{0},v_{0},u^{\prime}_{0},v^{\prime}_{0}):=w_{0}+W_{0}(t;\Delta_{a},u_{0},u^{\prime}_{0})+W_{0}(t;\Delta_{b},v_{0},v^{\prime}_{0}).$ (48)
@@ Appendix C Rainfall Hyetograph: triangular model
The graph of the instantaneous rain intensity $r(t)$
$r(t)=\left\{\begin{array}[]{ll}i_{a}\displaystyle\frac{t}{t_{a}},&0\leq t\leq t_{a},\\ i_{a}\displaystyle\frac{1-t/T_{d}}{1-\gamma},&t_{a}<t\leq T_{d},\\ 0,&{\rm otherwise}\end{array}\right.$
over the time interval $[0,T_{d}]$ has the shape of a triangle with $T_{d}$ and $i_{a}$ as its base and height, respectively (see Fig. 18).